Calculate the directivity of an antenna for θE = 30°, θH = 60


1
Expert's answer
2021-01-10T18:30:07-0500

"D = \frac{41257}{θ_E \times θ_H} \\

= \frac{41257}{30 \times 60} \\

= 23"

Answer: 23